A regulation size basketball has a volume of 455.9 cubic inches. Use this info to answer the following question. Approximate your answers to the nearest tenth.

1. The formula for the volume of a sphere is V=4/3pieR^3. What is the diameter of the basketball?

2. The formula for the circumference of the basketball is C=2pieR. If the circumference of a basketball is 29 inches, is it a regulation basketball?

Thanks.

Answer by nerdybill(5404)

A regulation size basketball has a volume of 455.9 cubic inches. Use this info to answer the following question. Approximate your answers to the nearest tenth.
.
1. The formula for the volume of a sphere is V=4/3pieR^3. What is the diameter of the basketball?
.
The problem states that the volume is 455.9 cubic inches and you know pi is 3.14 so plug it all into the formula and solve for R:
V=4/3pieR^3
455.9 = 4/3(3.14)R^3
1367.7 = 4(3.14)R^3
341.925 = (3.14)R^3
108.893 = R^3
4.8 inches = R
.
Diameter = 2R = 2(4.8)= 9.6 inches
.
2. The formula for the circumference of the basketball is C=2pieR. If the circumference of a basketball is 29 inches, is it a regulation basketball?
C=2pieR
29=2(3.14)R
14.5 = (3.14)R
4.6 inches = R
.
From, problem #1 though, we now know that a regulation size ball has a radius of 4.8 inches. Therefore, THIS is NOT a regulation basketball.